OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative

OpenStudy (amistre64):

and your derivative is?

OpenStudy (anonymous):

The derivative is 0? @amistre64

OpenStudy (freckles):

s(t)=-2-6t s'(t) is not 0

OpenStudy (freckles):

\[\frac{d}{dt}ct=c \frac{d}{dt}t=c \frac{dt}{dt}=c(1)=c\]